Clemson Software is considering a new project whose data are shown below. The required equipment has a 3-year tax life, after which it will be worthless, and it will be depreciated by the straight-line method over 3 years. Revenues and other operating costs are expected to be constant over the project's 3-year life. What is the project's Year 1 cash flow? Do not round the intermediate calculations and round the final answer to the nearest whole number. Equipment cost (depreciable basis) $97,000 Straight-line depreciation rate 33.333% Sales revenues, each year $60,000 Operating costs (excl. depr.) $25,000 Tax rate 35.0% ​ a. $29,979 b. $31,682 c. $36,792 d. $34,067 e. $30,660
Answer #1

Annual depreciation = 0.33333 * 97,000 = 32,333.01

Year 1 cash flow = (Sales - costs - depreciation)(1 - tax) + depreciation

Year 1 cash flow = (60,000 - 25,000 - 32,333.01)(1 - 0.35) + 32,333.01

Year 1 cash flow = 1,733.5435 + 32,333.01

Year 1 cash flow = $34,067